I have several homework problems on probablity theory, but I am unsure on how to begin. Is there a set formula that I could use for them all?
Here is my first problem:
A card is drawn at random from a standard 52-card deck. Find the probability that the card is not a queen.

Answer by edjones(8007)

52-4=48 non queens
48/52=12/13 the probability that the card is not a queen.
.
Ed